# MNSC-0001-2024.00

### v1

@@
[…]
MNSC-0001-2024.00
[…]
Jiang
[…]
AIGC-Chain: A Blockchain-Enabled Full Lifecycle Recording System for AIGC Product Copyright Management
[…]
AIGC-Chain: A Blockchain-Enabled Full Lifecycle Recording System for AIGC Product Copyright Management
[…]
As artificial intelligence technology becomes increasingly prevalent, Artificial Intelligence Generated Content (AIGC) is being adopted across various sectors. Although AIGC is playing an increasingly significant role in business and culture, questions surrounding its copyright have sparked widespread debate. The current legal framework for copyright and intellectual property is grounded in the concept of human authorship, but in the creation of AIGC, human creators primarily provide conceptual ideas, with AI independently responsible for the expressive elements. This disconnect creates complexity and difficulty in determining copyright ownership under existing laws. Consequently, it is imperative to reassess the intellectual contributions of all parties involved in the creation of AIGC to ensure a fair allocation of copyright ownership. To address this challenge, we introduce AIGC-Chain, a blockchain-enabled full lifecycle recording system designed to manage the copyright of AIGC products. It is engineered to meticulously document the entire lifecycle of AIGC products, providing a transparent and dependable platform for copyright management. Furthermore, we propose a copyright tracing method based on an Indistinguishable Bloom Filter, named IBFT, which enhances the efficiency of blockchain transaction queries and significantly reduces the risk of fraudulent copyright claims for AIGC products. In this way, auditors can analyze the copyright of AIGC products by reviewing all relevant information retrieved from the blockchain.
[…]
This research was supported by [grant number, funding agency].
[…]
@@ 1 Introduction
The widespread adoption of technologies such as AI painting and intelligent chat software like ChatGPT [^1] https://openai.com/chatgpt/ has drawn increased attention to Artificial Intelligence-Generated Content (AIGC) and has led to its broad application across various domains (Luo et al. 2022, Sun et al. 2022, Kandlhofer et al. 2016). AIGC leverages artificial intelligence techniques, including Generative Adversarial Networks (GAN) (Krichen 2023) and diffusion models, to train on existing datasets and generate new content based on specific input prompts. In 2018, an artwork created by artificial intelligence was sold at Christie’s auction house for \$432,500222https://edition.cnn.com/style/article/obvious-ai-art-christies-auction-smart-creativity/index.html, marking the first-ever sale of AIGC product worldwide. The price achieved underscored the commercial and cultural significance of AIGC. Nevertheless, this auction sparked significant debate regarding the copyright ownership of AIGC, primarily focusing on whether AIGC should be deemed copyrightable and, if it is, who should hold the copyright (He 2019).
The foundation of copyright and intellectual property law is built on the concept of the author as a human being. However, with the emergence of artificial intelligence, and its increasing role as a human assistant in the creation of various forms of work, this established legislative policy has been severely undermined. Consequently, the question of who should be recognized as the author of such works has become a highly controversial issue (GAFFAR and ALBARASHDI 2024).
In artificial intelligence generation, model users set core prompts such as keywords and select an appropriate AI model, collectively forming generative content (Chen et al. 2023b). AI models do not possess independent thought as a tool to replace human expression, while the keywords provided by the user are almost entirely conceptual and lack expressive form. Moreover, AI models are the crystallization of wisdom developers train based on a large amount of data. Consequently, it is crucial to assess the extent of knowledge contributions from various roles throughout the generation process of an AIGC product to ascertain copyright ownership. This process cannot be accomplished through mechanized procedural settings. Determining copyright ownership necessitates engaging third-party institutions and thoroughly evaluating the intellectual input and content contributions at different phases of the training process.
@@
However, the unpredictability of AI models renders it impossible to produce identical AIGC products even when utilizing the same models and prompts. Consequently, even if an individual genuinely participated in the creation of the AIGC product and possesses evidence to that effect, the irreproducibility of the outcome makes it challenging to verify the accuracy of the evidence provided. Attackers might still fabricate evidence to claim ownership. Therefore, meticulously recording every step of the AI-generated content creation process is vital to ensuring the transparency and traceability of AIGC products. In the event of copyright disputes, these accurate and unalterable records are crucial for a fair determination of copyright ownership.
Digital watermarking is a widespread technique utilized for digital copyright protection (Cox et al. 1997, Katzenbeisser et al. 2008, Celik et al. 2008, Li et al. 2023, Ren et al. 2024, Kamran et al. 2013). This technique involves embedding a unique identifier within the product, which serves as a means to identify the user and can be used to trace the infringing party in cases of copyright violation. The copyright of products protected by watermarking is generally clear and free from dispute. In the case of AIGC products, however, the attribution of copyright is often ambiguous. Watermarking facilitates the attribution of ownership during instances of copyright infringement. In stark contrast, the copyright ownership of AIGC products remains controversial.
Blockchain, with its inherent immutability and decentralized nature, has emerged as the preferred choice for establishing a comprehensive copyright management system for AIGC throughout its entire lifecycle. On the blockchain platform, every operation and data associated with the AIGC product can be meticulously recorded, with their integrity and chronological order verified through timestamps. With the implementation of this approach, AIGC copyright management has become more streamlined, enabling auditors to retrieve relevant information from the blockchain directly (Franke et al. 2024).
So far, blockchain technology has achieved notable success in copyright protection (Liang et al. 2021, Chen et al. 2023c, Ansori et al. 2023, Wang et al. 2022, Natgunanathan et al. 2022, Meng et al. 2018). The copyright ownership of products protected by blockchain, similar to that of digital watermarking, is also clearly identified. In other words, blockchain technology is primarily employed to enhance the traceability of copyright-related information. Hence, the conventional blockchain-enabled copyright protection approaches are not suitable for the copyright management of AIGC products. Although current research is exploring blockchain technology for managing copyrights of AIGC products (Liu et al. 2024a, b), these studies primarily focus on the recording of information directly related to the products, failing to adequately address the data management needs across the various stages of the lifecycle of an AIGC product.
@@
In this paper, we propose AIGC-Chain, a blockchain-enabled full lifecycle recording system for managing the copyright of AIGC products. As shown in Fig. 1, we leverage the blockchain to record all data produced throughout the lifecycle of the AIGC product, establishing a solid foundation of evidence to be utilized in the event of copyright disputes. In instances of copyright ownership disputes about AIGC products, the implementation of blockchain provides a reliable and tamper-proof evidentiary framework. This enables the direct retrieval of precise data from the blockchain, thereby assisting in the determination of the copyright holder for AIGC products.
Additionally, we present IBFT, an efficient copyright tracing method based on the Indistinguishable Bloom Filter (Li and Liu 2017), which accelerates on-chain transaction retrieval, significantly improving the efficiency of copyright management. Furthermore, we maintain a world state with eight attributes for each AIGC product, ensuring that intermediate data from the AIGC generation process is securely stored on the blockchain in an immutable form. This approach guarantees traceability, verifiability, and integrity for AIGC products. Our principal contributions can be encapsulated as follows:
  • We design a blockchain-enabled full lifecycle recording system for managing the copyright of AIGC products. It records the entire journey of an AIGC product, from initiating the content generation request to completing copyright transaction processes, providing robust data management capabilities for copyrights of AIGC products.
  • We introduce an efficient copyright tracing method, named IBFT, which is engineered to boost the efficiency of blockchain transaction retrieval. It ensures swift and dependable transaction validation, effectively reducing the risks associated with copyright fraud for AIGC items.
  • We have meticulously outlined the five steps within the full lifecycle recording of the AIGC product, thereby ensuring the effective management of copyrights for AIGC products.
@@ 2.1 The Copyright of AIGC
AIGC encompasses novels, articles, images, music, and other forms of content that are produced by artificial intelligence technology (Wang et al. 2023a). In recent years, the popularity of AI in creating artistic works has surged to unprecedented levels. This rapid expansion has precipitated numerous intricate challenges at the confluence of human and machine creativity, attracting considerable focus within the realm of copyright law (Burylo 2022). These developments have engendered several key issues as follows.
Copyright Ownership: Given that AIGC is produced by AI programs rather than human creators, its copyright has become a subject of debate. Some contend that the copyright of AIGC should be assigned to the developers or proprietors of the AI programs, whereas others argue that AIGC should be treated as a public resource, accessible for unrestricted use by anyone. GAFFAR et al. (GAFFAR and ALBARASHDI 2024) discussed the issue of copyright ownership in AIGC from a legal perspective, and attempted to clarify the definition of copyright and the methods for determining its ownership in AIGC. Several countries, including the United Kingdom, Ireland, New Zealand, and India, have proactively initiated legal reforms to more effectively address copyright challenges related to works that incorporate computer-generated elements and those created with the assistance of artificial intelligence (Wan and Lu 2021). However, these statutes fail to thoroughly address situations where a computer generates work with little or no direct human input, which could result in a lacuna in the legal framework.
Copyright Infringement: The generation process of AIGC products may involve operations such as copying, altering, and transforming existing works, which may potentially lead to copyright infringement (Mochram et al. 2022). Wang et al. (Wang et al. 2023b) meticulously summarized the security and privacy challenges inherent in AIGC, offering fresh insights for the research of AIGC. For instance, some images generated by AI may use existing images as training data without obtaining authorization from the original author, thereby triggering infringement disputes. Besides, attackers may leverage the assistance of bots to disseminate a large volume of information, republishing the work while pretending it is the original one (Liu et al. 2024a). Consumers are likely to first encounter the manipulated work provided by the attacker, thereby successfully altering the copyright ownership of the target work.
@@ 2.2 Blockchain-enabled Copyright Management
As a distributed ledger technology, blockchain has played a crucial role in the field of copyright protection, owing to its decentralized, tamper-proof, and transparent attributes (Zhang et al. 2023). Digital assets on the blockchain, such as NFTs, can be permanently recorded and verified, with their uniqueness and verifiability allowing them to flow between any network entities and generate value (Biais et al. 2023). Iyengar et al. (Iyengar et al. 2023) performed a thorough examination to assess the advantages and difficulties of applying blockchain technology in supply chain management and associated fields. Savilev (Savelyev 2018) identified several critical issues in copyright management, such as copyright transparency, control over digital copies, and licensing, and suggests that blockchain technology can effectively address these challenges.
@@
Various methods for copyright protection leveraging blockchain technology have been proposed thus far (Xiao et al. 2023b, Liang et al. 2021, Natgunanathan et al. 2022, Ansori et al. 2023, Sheng et al. 2020, Qureshi and Megías 2019). Wang et al. (Wang et al. 2022) had harnessed the strengths of zero-watermarking algorithms and blockchain to safeguard the copyright of images. Xiao et al. (Xiao et al. 2023a) proposed a fair payment scheme based on blockchain that has proven particularly effective in ensuring the secure exchange of copyrighted content. Chen et al. (Chen et al. 2023c) introduced a blockchain-based proactive defense strategy that offers the benefits of streamlined copyright transfers and preemptive penalties for infringement. Additionally, Xiao et al. (Xiao et al. 2023b) have developed a blockchain-based media-sharing network equipped with asymmetric fingerprinting technology, which meticulously records the processes of media content sharing and trading, thereby facilitating secure and dependable media exchange. Nonetheless, the fundamental question in AIGC copyright discussions is about the ownership of AIGC products. In contrast, the copyrights of products mentioned above are clear, serving to protect the rights of the identified copyright owners.
@@
Currently, research on blockchain-enabled copyright protection technologies for AIGC is still limited. Chen et al. (Chen et al. 2023a) discussed how blockchain provides a secure and reliable platform for AIGC. Nevertheless, he has not introduced a specific blockchain-based solution that effectively addresses the copyright concerns in AIGC. Oben AI published AIGC chain [^4] https://www.aigcchain.io/about in 2022, which allows users to contribute resources for training distributed AIGC models and acquiring rewards. However, the establishment of the entire ecosystem is still in its infancy, primarily aimed at providing crowdsourcing platforms for training AIGC models, without addressing issues related to AIGC copyright protection and commerce. Liu et al. (Liu et al. 2024b) designed a digital assets trading method based on price game and blockchain. Liu et al. (Liu et al. 2024a) leveraged blockchain to provide management services for AIGC products in cloud networks. However, these studies have not yet addressed how to manage the data generated throughout the entire lifecycle of an AIGC product, from its creation onwards.
@@ 3.1 Consortium Blockchain
A consortium chain is a blockchain network controlled by a limited number of organizations, known as consortium members. These members agree on the rules and protocols governing the network and work together to validate transactions and maintain the integrity of the ledger (Kim et al. 2022). Unlike public blockchains, where anyone can join and participate, consortium chains require explicit invitation and membership.
@@ 3.2 Indistinguishable Bloom Filter
Indistinguishable Bloom filter (IBF) (Li and Liu 2017) obfuscates data by simulating future unknown queries. As an extension of the Bloom Filter (BFs) (Tarkoma et al. 2012), IBF ensures efficient and accurate data retrieval while safeguarding data privacy. As illustrated in Fig. 3, IBF is constructed from two $m$-bit BFs placed side by side, where each pair of cells occupying the same positions in both BFs is referred to as a twin. Within each twin, one cell is designated as chosen and the other as unchosen, with the unchosen cell potentially holding either 0 or 1. IBF encompasses three primary algorithms (Liang et al. 2023): IBF.Setup(), IBF.Insert(), and IBF.Check().
  • IBF.Setup$(1^{\eta},k,m)\to\left\{H_{key},H,IBF\right\}$: Construct $k$ pseudo-random hash functions $H_{key}=\left\{h_{1},h_{2},\cdots,h_{k+1}\right\}$ using the keyed-hash message authentication code (HMAC) and a hash function $H(x)=\mathbb{H}\%2$, where $\mathbb{H}$ is a pseudorandom hash function. Besides, generate two $m$-bit sets as $IBF$.
  • IBF.Insert$(IBF,v,H_{key},H,\gamma)\to IBF$: For each value $v$, set $IBF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1$ and $IBF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0$, where $i\in\left[1,k\right]$ and $\gamma$ is a random number.

### v2

# Trustworthy AIGC Copyright Management with Full Lifecycle Recording and Multi-party Supervision in Blockchain

@@
Abstract
As artificial intelligence technology becomes increasingly widespread, AI-generated content (AIGC) is gradually penetrating into many fields. Although AIGC plays an increasingly prominent role in business and cultural communication, the issue of copyright has also triggered widespread social discussion. The current legal system for copyright is built around human creators, yet in the realm of AIGC, the role of humans in content creation has diminished, with the creative expression primarily reliant on artificial intelligence. This discrepancy has led to numerous complexities and challenges in determining the copyright ownership of AIGC within the established legal boundaries. In view of this, it is necessary to meticulously record contributions of all entities involved in the generation of AIGC to achieve a fair distribution of copyright. For this purpose, this study thoroughly records the intermediate data generated throughout the full lifecycle of AIGC and deposits them into a decentralized blockchain system for secure multi-party supervision, thereby constructing a trustworthy AIGC copyright management system. In the event of copyright disputes, auditors can retrieve valuable proof from the blockchain, accurately defining the copyright ownership of AIGC products. Both theoretical and experimental analyses confirm that this scheme shows exceptional performance and security in the management of AIGC copyrights.
Index Terms:
@@ 1 Introduction
The widespread adoption of technologies such as AI painting and intelligent chat software like ChatGPT [^1] https://openai.com/chatgpt/ has drawn increased attention to Artificial Intelligence-Generated Content (AIGC) and has led to its broad application across various domains[32, 36, 37]. AIGC leverages artificial intelligence techniques, including Generative Adversarial Networks (GAN) [30] and diffusion models, to train on existing datasets and generate new content based on specific input prompts. In 2018, an artwork created by artificial intelligence was sold at Christie’s auction house for \$432,500222https://edition.cnn.com/style/article/obvious-ai-art-christies-auction-smart-creativity/index.html, marking the first-ever sale of AIGC product worldwide. The price achieved underscored the commercial and cultural significance of AIGC. However, this auction sparked a significant debate on AIGC copyright ownership, involving the distribution of rights among model providers, users, and other possible creative contributors of AIGC [18].
Copyright refers to the legal rights of creators to their literary, artistic and scientific works, with the aim of protecting the creative intellectual output of creators from unauthorized use by others, and encouraging innovation and cultural development. Traditional copyright law assumes that works are created by human creators. while the creation of AIGC is the result of collaboration among multiple participants. For instance, model providers develop AI models through extensive data training and computation, while users contribute to the content generation process by supplying prompts based on their personal experience and needs, which are then used to operate the AI content generation through the model [38]. The role of humans in content creation has diminished, raising new questions about whether the copyright of AIGC belongs to the AI technology itself, model providers, users, or other relevant parties. This challenges the requirement of human author in traditional copyright law. Moreover, copyright law generally necessitates that a work demonstrates a degree of originality, yet determining whether AI-generated content satisfies this criterion and assessing the originality of such content pose complex and undefined challenges [19]. As a result, AIGC frequently encounters quandaries related to copyright infringement.
@@
However, traditional data storage methods, with their centralized architecture, are vulnerable to attacks and tampering [34]. It allows each participant in the process of AIGC to potentially submit altered or fabricated proof that favors their interests during copyright attribution, affecting the determination of copyright ownership for AIGC. To address this flaw, there is an urgent need for a more secure and transparent recording system that ensures the integrity of proof. In this context, the application of blockchain appears to be crucial. Blockchain, with its unique decentralized structure and advanced encryption algorithms, can provide AIGC with a highly robust, decentralized, and manageable end-to-end copyright management platform [27]. By uploading the details of AI-generated content creation, copyright declarations, and information on all critical steps to the blockchain, these records are secured against unauthorized alterations once recorded [4]. Moreover, the timestamp of blockchain assigns an exact temporal verification to each record, enhancing the integrity and credibility of the proof while also ensuring the immutability of the transaction timeline, thus providing strong technical support for the protection and legal defense of AIGC copyrights. The transparency and traceability of blockchain provide convenience for copyright regulation, making the oversight process more open and fair, significantly improving the trustworthiness of the entire AIGC ecosystem.
So far, significant advancements have been achieved in the application of blockchain for copyright protection [8, 9, 11, 12, 23, 29], particularly in enhancing the security and traceability of copyright information storage. However, the majority of these applications are tailored to scenarios where the attribution of copyrights is unambiguous. The copyright of each AIGC product involves the contributions of multiple participants throughout various stages of content generation, which are often difficult to delineate and quantify. Evidently, the existing blockchain-enabled copyright protection strategies have not adequately addressed this complexity. Currently, research efforts are underway to apply blockchain to the management of AIGC copyrights [7, 33]. Nevertheless, these studies typically operate on the assumption that the copyright of AIGC is owned by a single entity, with a primary focus on the copyright registration and transaction management of the product itself. These research endeavors have not yet comprehensively considered the contributions of various participants throughout the content generation and their implications on copyright throughout the full lifecycle of AIGC products.
@@
In this paper, we facilitate trustworthy AIGC copyright management by meticulously recording intermediate data generated throughout the full lifecycle of AIGC and leveraging blockchain to establish a decentralized multi-party supervision framework. As illustrated in Fig. 1, we pre-recorded the intermediate data generated at each stage of the AIGC lifecycle on the AIGC-Chain. In the instance of copyright disputes, we draw upon the data from the decentralized ledger to provide proof, thus offering a solid foundation and support for resolving copyright conflicts. Specifically, we maintain a world state with eight attributes for each AIGC product, ensuring that intermediate data of the generation process is securely stored on the blockchain in an immutable manner. This methodology establishes a robust framework ensuring traceability, verifiability, and integrity of AIGC products, effectively overcoming the constraints inherent in conventional content generation paradigms. By moving beyond both the anthropocentric authorship model and reductive instrumental approaches, it creates a reliable infrastructure for synergistic human-AI co-creation. Our main contributions can be summarized as follows:
  • We introduce AIGC-Chain, a trustworthy copyright management framework for the full lifecycle of AIGC, meticulously designing five transaction phases required for documentation and details of their state transition on the blockchain, aiming to achieve holistic management and complete recording of the status of AIGC.
  • We establish a decentralized multi-party supervision framework with blockchain, which enables automated copyright proof collection. This framework is augmented by IPFS technology to safeguard the the security and reliable storage of AIGC copyright.
  • We develop CTrace, a copyright tracing method based on bloom filter, which greatly improves the efficiency and accuracy of transaction query and verification in blockchain, provides strong proof support for copyright disputes, and thus realizes trustworthy AIGC copyright management.
@@ 2.1 The Copyright of AIGC
AIGC encompasses novels, articles, images, music, and other forms of content that are produced by artificial intelligence technology[28]. In recent years, the popularity of AI in creating artistic works has surged to unprecedented levels. This rapid expansion has precipitated numerous intricate challenges at the confluence of human and machine creativity, attracting considerable focus within the realm of copyright law[41]. These developments have engendered several key issues as follows.
Copyright Ownership: Given that AIGC is produced by AI programs rather than human creators, its copyright has become a subject of debate. Some contend that the copyright of AIGC should be assigned to the developers or proprietors of the AI programs, whereas others argue that AIGC should be treated as a public resource, accessible for unrestricted use by anyone. GAFFAR et al.[19] discussed the issue of copyright ownership in AIGC from a legal perspective, and attempted to clarify the definition of copyright and the methods for determining its ownership in AIGC. Several countries, including the United Kingdom, Ireland, New Zealand, and India, have proactively initiated legal reforms to more effectively address copyright challenges related to works that incorporate computer-generated elements and those created with the assistance of artificial intelligence[20]. However, these statutes fail to thoroughly address situations where a computer generates work with little or no direct human input, which could result in a lacuna in the legal framework.
Copyright Infringement: The generation process of AIGC products may involve operations such as copying, altering, and transforming existing works, which may potentially lead to copyright infringement[21]. Wang et al.[26] meticulously summarized the security and privacy challenges inherent in AIGC, offering fresh insights for the research of AIGC. For instance, some images generated by AI may use existing images as training data without obtaining authorization from the original author, thereby triggering infringement disputes. Besides, attackers may leverage the assistance of bots to disseminate a large volume of information, republishing the work while pretending it is the original one[7]. Consumers are likely to first encounter the manipulated work provided by the attacker, thereby successfully altering the copyright ownership of the target work.
@@ 2.2 Blockchain-enabled Copyright Management
As a distributed ledger technology, blockchain has played a crucial role in the field of copyright protection, owing to its decentralized, tamper-proof, and transparent attributes[6]. Digital assets on the blockchain, such as NFTs, can be permanently recorded and verified, with their uniqueness and verifiability allowing them to flow between any network entities and generate value[5]. Iyengar et al. [3] performed a thorough examination to assess the advantages and difficulties of applying blockchain technology in supply chain management and associated fields. Savilev [24] identified several critical issues in copyright management, such as copyright transparency, control over digital copies, and licensing, and suggests that blockchain technology can effectively address these challenges.
@@
Various methods for copyright protection leveraging blockchain technology have been proposed thus far[22, 8, 23, 11, 39, 40]. Wang et al. [12] had harnessed the strengths of zero-watermarking algorithms and blockchain to safeguard the copyright of images. Xiao et al.[10] proposed a fair payment scheme based on blockchain that has proven particularly effective in ensuring the secure exchange of copyrighted content. Chen et al.[9] introduced a blockchain-based proactive defense strategy that offers the benefits of streamlined copyright transfers and preemptive penalties for infringement. Additionally, Xiao et al.[22] have developed a blockchain-based media-sharing network equipped with asymmetric fingerprinting technology, which meticulously records the processes of media content sharing and trading, thereby facilitating secure and dependable media exchange. Nonetheless, the fundamental question in AIGC copyright discussions is about the ownership of AIGC products. In contrast, the copyrights of products mentioned above are clear, serving to protect the rights of the identified copyright owners.
@@
Currently, research on blockchain-enabled copyright protection technologies for AIGC is still limited. Chen et al. [25] discussed how blockchain provides a secure and reliable platform for AIGC. Nevertheless, he has not introduced a specific blockchain-based solution that effectively addresses the copyright concerns in AIGC. OpenAI published AIGC chain [^4] https://www.aigcchain.io/about in 2022, which allows users to contribute resources for training distributed AIGC models and acquiring rewards. However, the establishment of the entire ecosystem is still in its infancy, primarily aimed at providing crowdsourcing platforms for training AIGC models, without addressing issues related to AIGC copyright protection and commerce. Liu et al.[33] designed a digital assets trading method based on price game and blockchain. Liu et al. [7] leveraged blockchain to provide management services for AIGC products in cloud networks. However, these studies have not yet addressed how to manage the data generated throughout the entire lifecycle of an AIGC product, from its creation onwards.
@@ 3.1 Consortium Blockchain
A consortium chain is a blockchain network controlled by a limited number of organizations, known as consortium members. These members agree on the rules and protocols governing the network and work together to validate transactions and maintain the integrity of the ledger[17]. Unlike public blockchains, where anyone can join and participate, consortium chains require explicit invitation and membership.
@@ 3.2 Indistinguishable Bloom Filter
Indistinguishable Bloom filter (IBF) [16] obfuscates data by simulating future unknown queries. As an extension of the Bloom Filter (BFs) [14], IBF ensures efficient and accurate data retrieval while safeguarding data privacy. As illustrated in Fig. 3, IBF is constructed from an $m$-bit Twin Bloom Filter[15]. Within each twin, one cell is designated as chosen and the other as unchosen, with the unchosen cell potentially holding either 0 or 1. IBF encompasses three primary algorithms: IBF.Setup, IBF.Initi, IBF.Insert, and IBF.Check.
  • IBF.Setup$(1^{\eta},k,m)\to\left\{H_{key},H\right\}$: Construct $k$ pseudo-random hash functions $H_{key}=\left\{h_{1},h_{2},\cdots,h_{k+1}\right\}$ using the keyed-hash message authentication code (HMAC) and a hash function $H(x)=\mathbb{H}\%2$, where $\eta$ is the security parameter, $k$ is the number of hash functions, and $\mathbb{H}$ is a pseudorandom hash function.
  • IBF.Initi($m,H_{key},H,\gamma\to IBF$) Calculate $H(h_{k+1}(j)\oplus\gamma)$ to determine which cell in each twin is chosen, where $\gamma$ is a random number and $j\in\{0,m-1\}$.
  • IBF.Insert$(IBF,v,H_{key},H,\gamma)\to IBF$: For each value $v\in V$, set $IBF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1$ and $IBF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0$, where $i\in\left[1,k\right]$.
@@
The correctness of IBF is established if, for all parameters $\eta$, $m$, $k$, $\gamma$, and any value $v$, following the execution of IBF.Setup, IBF.Init, IBF.Insert, and IBF.Check, the function IBF.Check returns a value of 1 when $v^{\prime}=v$, otherwise
$Pr[IBF.Check(IBF,v^{{}^{\prime}},\gamma)]=0]=1-\textnormal{negl}_{\textnormal{IBF}}(v,k,|V|),$
where the false positive rate of IBF is identical to that of traditional BF, i.e., neglIBF($m,k,|V|$) = neglBF($m,k,|V|$) $\approx(1-(1-\frac{1}{m})^{k|V|})^{k}\approx(1-e^{\frac{-k|V|}{m}})^{k}$.